OpenStudy (anonymous):

You plant 40 pots with seedlings. Eight of the pots contain tomato plants. What percent of your seedlings are not tomato plants?

OpenStudy (kropot72):

The number of pots tha do not contain tomato plants = 40 - 8 The total number of pots with seedlings = 40 The percentage of pots that do not contain tomato plants is: \[\frac{40-8}{40}\times 100\]